Supposed Ghana Standards Boards collected 40 electric bulbs from ELECTRO-NETS LTD for testing and 10 of them happened to be defective. If 4 of the bulbs were randomly selected one after the other, what is the probability that all the 4 bulbs are defective assumed sampling without replacements.


1
Expert's answer
2021-07-07T13:59:54-0400

Let "X=" the number of defective bulbs among 4 selected


"P(X=4)=\dfrac{10}{40}(\dfrac{9}{39})(\dfrac{8}{38})(\dfrac{7}{37})=\dfrac{21}{9139}"

"\approx0.0023"